Please show work.
Consider the functions:
f(x)=x^2+4
g(x)=x
a. Find f(g(x))
b. f(g(1))

Answer by fractalier(6550)

Since g(x) = x, we have
f(g(x)) = f(x) = x^2 + 4
f(g(1)) = 1^2 + 4 = 5